OpenStudy (anonymous):

Find the general solution of the differential equation 3ln(y)y′−ty=0

OpenStudy (anonymous):

last one!

OpenStudy (anonymous):

do you happen to know the integrating factor method?

OpenStudy (anonymous):

unless it has a different name, no.

OpenStudy (zarkon):

this is a separable diffeq...that is easy to integrate

OpenStudy (anonymous):

so far i'm at \[\frac{ln(y)}{y}dy=\frac{t}{3}dt\]

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

zarkon is right, use u substitution on ln(y)

OpenStudy (anonymous):

\[\frac{ln(y)^2}{2}=\frac{t^2}{6}\]

OpenStudy (zarkon):

+c

OpenStudy (anonymous):

oh ty

OpenStudy (anonymous):

awesome :)

OpenStudy (anonymous):

\[ln(2y)=\frac{t^2}{3}+C\]

OpenStudy (zarkon):

?

OpenStudy (anonymous):

edit: \[2ln(y)\]

OpenStudy (zarkon):

no

OpenStudy (zarkon):

only if it was \[\ln(y^2)=2\ln(y)\] but it is not

OpenStudy (anonymous):

ah

OpenStudy (zarkon):

just take the sqrt of both sides

OpenStudy (anonymous):

\[ln(y)=(sqrt{\frac{t^2}{3}+C})\]

OpenStudy (anonymous):

sqrt syntax?

OpenStudy (zarkon):

\sqrt{}

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is getting harder and harder to type lol \[y = e^{\sqrt{\frac{t^2}{3}+C}}\]

OpenStudy (anonymous):

now you see why i like to write it out on paper lolol

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

Hes a catch. The general solution is different if y is initially greater than or less than 1. what does that mean?

OpenStudy (zarkon):

when you took the sqrt...you introduced two solution + and - so it was possible the ln was negative

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

is it pos/neg on the sqrt?

OpenStudy (zarkon):

yes..depending on the initial condition

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

bingo.

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

Thanks guys! good night.

OpenStudy (anonymous):

lates :P

OpenStudy (anonymous):

do you have the answer from the book?